OpenStudy (anonymous):

Sue buys 5 notebooks and 3 binders for 21.56. Tim buys 4 notebooks and 2 binders for 15.80. how much is each item? How do I solve this??

OpenStudy (anonymous):

let x= # of notebooks let y= # of binders \[5x+3y=21.56\] \[4x+2y=15.8\] Now do you know how to solve it?

OpenStudy (anonymous):

I get that same set up but the next step is what I don't understand. I am told to multiply the top row by 4...4(5x+3y)=21.56 and the next row by -5(4x+2y)=15.80....why am I supposed to multiply by the 4 and -5?

OpenStudy (anonymous):

Ah yes, you see, there are different ways to solve that, you are doing the elimination method. You multiply the top row by 4 to make 20x, and what ever the rest of the equation is multiplied by 4. You multiply the bottom row by -5 to get -20x, and whatever the rest of the equation is multiplied by -5. Then you add the 2 equations, so: \[4(5x+3y)=21.56\] \[=20x+12y=21.56\] \[-5(4x+2y)=15.80\] \[=-20x-10y=15.80\] Then you add them: 20x+-20x=0 12y+10y=2y 21.56+15.80= 37.36 So: \[2y=37.36\] \[y=18.68\] Then you plug that in to one of the original equations to find x

OpenStudy (anonymous):

Sorry, I meant -12y+10y=2y

OpenStudy (anonymous):

I was doing the elimination method and didn't know it. Thanks for the help, I really appreciate the help